Ca5(PO4)3F + 5H2SO4 = 3H3PO4 + 5CaSO4 + HF if 24.5 g Ca(PO4)3F is mixed with 89.7 g H2SO4, how much H3PO4 will form?



NEED IT TODAY

1
Expert's answer
2021-11-19T01:07:07-0500

The quantity of Ca5(PO4)3F:

(24.5 g) / (504.3 g/mol) = 0.0486 mol

The quantity of H2SO4:

(89.7 g) / (98.1 g/mol) = 0.9144 mol

The ratio of these reactants must be 1:5, so H2SO4 is in excess.

Therefore, the quantity of H3PO4 (according to balanced reaction equation):

3 * (0.0486 mol) = 0.1458 mol

The mass of H3PO4:

(0.1458 mol) * (98 g/mol) = 14.3 g

Answer: 14.3 g